# Efficient algorithms for collecting the statistics of large-scale IP address data

### v2

@@ Abstract
Compiling the statistics of large-scale IP address data is an essential task in network traffic measurement. The statistical results are used to evaluate the potential impact of user behaviors on network traffic. This requires algorithms that are capable of storing and retrieving a high volume of IP addresses within time and memory constraints. In this paper, we present two efficient algorithms for collecting the statistics of large-scale IP addresses that balance time efficiency and memory consumption. The proposed solutions take into account the sparse nature of the statistics of IP addresses while building the hash function and maintain a dynamic balance among layered memory blocks. There are two layers in the first proposed method, each of which contains a limited number of memory blocks. Each memory block contains 256 elements of size $256\times 8$ bytes for a 64-bit system. In contrast to built-in hash mapping functions, the proposed solution completely avoids expensive hash collisions while retaining the linear time complexity of hash-based solutions. Moreover, the mechanism dynamically determines the hash index length according to the range of IP addresses, and can balance the time and memory constraints. In addition, we propose an efficient parallel scheme to speed up the collection of statistics. The experimental results on several synthetic datasets show that the proposed method substantially outperforms the baselines with respect to time and memory space efficiency.
keywords:
large-scale IP addresses, memory blocks, hash table, sorting, network traffic
journal: Information Processing Letterst1t1footnotetext: This work was partially supported by National Natural Science Foundation of China (NSFC) under Grants 61303015 and Sichuan Science and Technology Program under Grant 2021YJ0078
@@ 1 Introduction
A number of statistical algorithms for solving this problem have been studied in the last few decades Jing05DLBS ; Kapur2013SA ; Klein2013Sorting ; Fredman2014Sorting ; Agapitos2016RSA . A classic divide-and-conquer strategy has been proposed in which IP addresses are first divided into multiple subsets. Then, each subset is individually computed using a statistics collection method, such as a top $k$ trie algorithm Jing2013SEDS . Finally, to merge and sort the results of the multiple subsets, sorting algorithms (e.g., bubble sort, insertion sort, merge sort, selection sort, or quick sort) are used Cormen2001IA ; Hammad2015VSA . However, the reduction of computational cost is an intractable problem in the sorting procedure Puglisi2005SSA ; Louza2017Sorting ; Abdel2017CFSA . For example, the average complexity and worst-case complexity of the bubble, insertion, and selection sort algorithms are $O\left({{n^{2}}}\right)$ Hammad2015VSA ; Shabaz2019SA , where $n$ represents the number of unsorted records. This indicates that the merging and sorting step of the multiple subsets occupies a large amount of memory and is extremely computationally costly to perform when collecting the statistics of millions or tens of millions of IP addresses. Therefore, statistics collection algorithms using the divide-and-conquer strategy still face several challenges caused by the rapid increase of the large-scale records, such as bounded memory and computational cost restrictions.
@@
A hash table is an effective method for collecting the statistics of IP addresses Sanders2015HS . It uses a hash function to compute a hash codes for an array of buckets with the statistical results. The hash function assigns each key to a unique bucket for each IP address. Unfortunately, the hash function can generate the same hash code for more than one IP address. With the increase in the generation of big data, millions or tens of millions of records have become ubiquitous in network traffic. Therefore, this approach could cause several hash collisions, especially for a large number of IP addresses. Although many strategies can be employed to avoid collisions, such as linear probing, quadratic probing, and double hashing, they require extra storage space and computation.
@@
The statistics collection algorithm should be stable, effective, and efficient for large-scale records. To overcome the disadvantages of general statistics collection methods, a number of parallel techniques have been developed for large-scale records by optimizing the efficiency and complexity. For example, these algorithms have been extended to the corresponding parallel structures for parallel hardware architectures, such as many-core and multi-core platforms Satish2009SA ; Singh2018GPU . However, these parallel statistics collection algorithms are still the parallel computing variants of the original algorithms, and they continue to ignore the importance of the original characteristics of the unsorted records.
@@ 2.1 Classical sort techniques
Bubble sort is a classical sorting algorithm in which each element in a list is compared with its neighboring elements and swapped until they are in the desired order Shutler2008SA . Bubble sort leads to $\left({n-1}\right)$ number of passes and $\frac{{n\left({n-1}\right)}}{2}$ number of iterations if $n$ elements are given. Insertion sort is a simple and efficient sorting algorithm that iteratively takes one element and finds its appropriate position in the sorted list by comparing it with neighboring elements. It becomes less efficient as the number of records increases. Selection sort determined the smallest number in an unsorted list and swaps it with the first number in the sorted list. Then, it finds the next smallest element from the remaining list and swaps with the second element in the sorted list. Consequently, the number of sorted elements at the top of the list increases while the rest remain unsorted. Merge sort, which is based on the divide-and-conquer principle, repeatedly divides the array into two halves and then combines them in a sorted manner For more details of classical sort algorithms, such as quick sort and heap sort, we refer the reader to the comments in Idrizi2017SA ; Cormen2001IA . It is well known that computational cost and required memory are the primary concerns in sorting algorithms Rusu2018Sorting ; Jukna2020Sorting . Table 1 shows the time complexity of the best case, average case, and worst case of several classical sorting algorithms Cormen2001IA ; Kocher2014SA ; Faujdar2015SA .
Table 1: Time complexity of various sorting algorithms.
@@ 2.2 Parallel hardware architectures
The hardware architecture of modern processors usually consists of more than two independent central processing units (CPUs) or graphics processing units (GPUs). Parallel software platforms can be implemented using high-level programming frameworks for specific hardware architectures Chen2009SA . The Compute Unified Device Architecture (CUDA) is a parallel computing platform for general computing on GPUs. Most parallel sorting algorithms are variants of standard, well-known sorting algorithms adapted to GPU hardware architecture. For example, Cederman designed a quick sort for the GPU platform Cederman2008 , and Peters proposed an adaptive bitonic sorting algorithm with a bitonic tree for GPUs Peters2011BS . The parallel computation of sorting algorithms is considered to be the most efficient way of sorting elements on parallel hardware architectures Singh2018GPU .
@@ 3.1 Problem Formulation
Assume that IP flow data $FD$ is a sequence of IP records, that is, $FD=\left\{\left({p_{1},x_{1}}\right),\left({p_{2},x_{2}}\right),...,\left({p_{n},x_{n}}\right)\right\}\in{\mathbb{R}^{m\times n}}$ and ${n\geq 1{e^{6}}}$, where each pair of elements $\left({p_{i},x_{i}}\right)$ $\left({i\in\left[{1,n}\right]}\right)$ consists of an IP address $p_{i}$ and a set of corresponding user behavior attributes $x_{i}$. Given a finite set of IP addresses $P=\left\{{{p_{1}},{p_{2}},...,{p_{n}}}\right\}\in{\mathbb{R}^{n}}$, the purpose of the IP address statistics task is to efficiently determine the first $k$ IP addresses of the most frequent occurrences in $P$, where $k\ll n$.
@@ 3.3.1 First method: TLMB
We formally present a storage strategy for IP addresses that consists of two layers that consist of a limited number of memory blocks. The first layer contains $256\times 256$ memory blocks. The first three parts of the IP address can be mapped into the corresponding position of the element in a particular memory block of the first layer according to the individual values of the three parts. We allocate a memory block in the other layer for the IP address when its first three parts are initially given. Each element of a memory block in this layer stores the number of occurrences of the corresponding IP address. Figure 2 shows an example of the relationship mapping between the memory blocks of two layers and an IP addresses. The positions of the first three dimensionalities of the sparse matrix can be mapped to the elements of the memory blocks in the first layer. The positions of the final dimensionality of the sparse matrix can be indexed by combining the starting address of the memory block in the second layer with the integer value of the four parts of the IP address.
3.3.2 Second method: SSMB
@@ 3.5.2 Parallel computation mechanism of SSMB
Table 2: Statistics of the three synthetic datasets.
Data sets IP Records Individual IP Addresses Size
1 5, 000, 000 50, 000 77.5 MB
2 10, 000, 000 100, 000 155 MB
3 50, 000, 000 500, 000 775 MB
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
1 10 1458.14 (2.32) 15.18 (0.07) 2.51 (0.01) 18.43 (0.05)
100 1458.56 (2.76) 15.21 (0.04) 2.52 (0.01) 18.43 (0.06)
2 10 2927.23 (11.64) 17.41 (0.11) 4.92 (0.01) 27.59 (0.05)
100 2934.65 (28.93) 17.45 (0.05) 4.95 (0.01) 27.65 (0.06)
3 10 14547.09 (13.95) 35.33 (0.09) 24.22 (0.07) 101.01 (0.17)
100 14566.36 (26.78) 35.39 (0.05) 24.24 (0.04) 101.20 (0.2)
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
1 10 43.63 (0.12) 16332.81 (0.07) 189.61 (0.16) 139.77 (0.13)
100 43.73 (0.06) 16332.77 (0.05) 189.64 (0.07) 139.79 (0.07)
2 10 74.8 (0.12) 16332.75 (0.12) 239.11 (0.1) 139.74 (0.13)
100 74.83 (0.05) 16332.77 (0.07) 239.06 (0.15) 139.7 (0.12)
3 10 324.45 (0.99) 16332.59 (0.06) 628.33 (0.07) 139.6 (0.15)
100 324.4 (1.02) 16332.6 (0.16) 628.27 (0.01) 139.71 (0.16)
@@ 4.1 Experimental Settings
In this section, we evaluate the performance of the proposed methods  [^1] https://github.com/chenjie20/IPStatistics on three synthetic datasets that contain 5 million, 10 million, and 50 million randomly generated IP records. Each individual IP address contains one or more of IP records. The average number of IP records is 100 for each individual IP address. Parameter $k$ represents the number of frequently occurring IP addresses. The statistics of these datasets are summarized in Table 2.
@@ 2. $\bullet$
Two metrics were employed to evaluate the sorting performance, that is, computational cost and memory use. All experiments were implemented using the C language on a Windows platform with an Intel i7-9700k CPU and 32 GB RAM.

### v3

# Foshan 528137, P.R. China

@@ Abstract
Compiling the statistics of large-scale IP address data is an essential task in network traffic measurement. The statistical results are used to evaluate the potential impact of user behaviors on network traffic. This requires algorithms that are capable of storing and retrieving a high volume of IP addresses within time and memory constraints. In this paper, we present two efficient algorithms for collecting the statistics of large-scale IP addresses that balance time efficiency and memory consumption. The proposed solutions take into account the sparse nature of the statistics of IP addresses while maintaining a dynamic balance among layered memory blocks. There are two layers in the first proposed method, each of which contains a limited number of memory blocks. Each memory block contains 256 elements of size $256\times 8$ bytes for a 64-bit system. In contrast to built-in hash mapping functions, the proposed solution completely avoids expensive hash collisions while retaining the linear time complexity of hash-based solutions. Moreover, the mechanism dynamically determines the hash index length according to the range of IP addresses, and can balance the time and memory constraints. In addition, we propose an efficient parallel scheme to speed up the collection of statistics. The experimental results on several synthetic datasets show that the proposed method substantially outperforms the baselines with respect to time and memory space efficiency.
Keywords: large-scale IP addresses, memory blocks, hash table, sorting, network traffic.
@@ 1 Introduction
Each IP address serves as a unique identifier for a device on a network. Analyzing large-scale IP address data can reveal vulnerabilities in the network traffic measurements; this enables administrators to discover and resolve weak spots before they are exploited. As a result, potential cybersecurity risks can be effectively reduced in various network applications. Moreover, analyzing such data helps in understanding traffic patterns and behaviors in network traffic measurement process. The collection and analysis of large-scale IP address data provide valuable insights that can guide more reliable and efficient network systems; this helps administrators balance loads and improve the overall performance of the network. Therefore, collecting statistics from large-scale IP address data is an essential task for the efficient, secure, and scalable management of networks.
@@
A number of statistical algorithms for solving this problem have been studied in the last few decades [11], [15], [16], [8], [2]. A classic divide-and-conquer strategy has been proposed in which IP addresses are first divided into multiple subsets. Then, each subset is individually computed using a statistics collection method, such as a top $k$ trie algorithm [12]. Finally, to merge and sort the results of the multiple subsets, sorting algorithms (e.g., bubble sort, insertion sort, merge sort, selection sort, or quick sort) are used [9], [6]. However, the reduction of computational cost is an intractable problem in the sorting procedure [21], [18], [1]. For example, the average complexity and worst-case complexity of the bubble, insertion, and selection sort algorithms are $O\left({{n^{2}}}\right)$ [9], [24], where $n$ represents the number of unsorted records. This indicates that the merging and sorting step of the multiple subsets occupies a large amount of memory and is extremely computationally costly to perform when collecting the statistics of millions or tens of millions of IP addresses. Therefore, statistics collection algorithms using the divide-and-conquer strategy still face several challenges caused by the rapid increase of the large-scale records, such as bounded memory and computational cost restrictions.
@@
A hash table is an effective method for collecting the statistics of IP addresses [19]. It uses a hash function to compute a hash codes for an array of buckets with the statistical results. The hash function assigns each key to a unique bucket for each IP address. Unfortunately, the hash function can generate the same hash code for more than one IP address. With the increase in the generation of big data, millions or tens of millions of records have become ubiquitous in network traffic. Therefore, this approach could cause several hash collisions, especially for a large number of IP addresses. Although many strategies can be employed to avoid collisions, such as linear probing, quadratic probing, and double hashing, they require extra storage space and computation.
@@
The statistics collection algorithm should be stable, effective, and efficient for large-scale records. Recent advancements in sorting techniques have concentrated on improving the efficiency and scalability of algorithms for processing large-scale data.To overcome the disadvantages of general statistics collection methods, a number of parallel techniques have been developed for large-scale records by optimizing the efficiency and complexity. For example, these algorithms have been extended to the corresponding parallel structures for parallel hardware architectures, such as many-core and multi-core platforms [3], [27], [23],[26].
An IP address consists of a series of numbers separated by periods. However, the sorting techniques ignore the importance of the original characteristics of the unsorted IP records. Additionally, parallel statistics collection algorithms are simply parallel computing implementations of the original algorithms. The uniqueness of the IP addresses is vital for distinguishing between the different devices. An IPv4 address is typically represented in four parts. Consequently, collecting statistics from large-scale IP address data still face two significant challenges. First, the unique structure of unsorted IP records is ignored when large-scale IP address data are collected. Second, the extension of the collection task to the corresponding parallel computation paradigm deserves a further investigation into parallel hardware architectures.
@@ 2.1 Classical sort techniques
Bubble sort is a classical sorting algorithm in which each element in a list is compared with its neighboring elements and swapped until they are in the desired order [25]. Bubble sort leads to $\left({n-1}\right)$ number of passes and $\frac{{n\left({n-1}\right)}}{2}$ number of iterations if $n$ elements are given. Insertion sort is a simple and efficient sorting algorithm that iteratively takes one element and finds its appropriate position in the sorted list by comparing it with neighboring elements. It becomes less efficient as the number of records increases. Selection sort determined the smallest number in an unsorted list and swaps it with the first number in the sorted list. Then, it finds the next smallest element from the remaining list and swaps with the second element in the sorted list. Consequently, the number of sorted elements at the top of the list increases while the rest remain unsorted. Merge sort, which is based on the divide-and-conquer principle, repeatedly divides the array into two halves and then combines them in a sorted manner For more details of classical sort algorithms, such as quick sort and heap sort, we refer the reader to the comments in [10], [6]. It is well known that computational cost and required memory are the primary concerns in sorting algorithms [22], [14]. Table 1 shows the time complexity of the best case, average case, and worst case of several classical sorting algorithms [6], [17], [7]. These sorting algorithms can be used to find the first $k$ IP addresses of the most frequent occurrences from the statistical results obtained for IP address data.
Table 1: Time complexity of various sorting algorithms [6], [17], [7].
@@
2.2 Accelerating large-scale sorting techniques
Sorting is an essential part of modern computing. Significant efforts have recently been dedicated to accelerating large-scale sorting techniques [3], [13], [27]. For example, Alhabboub et al. improve the computation efficiency of the classical QuickSort algorithm by combining with parallel implementations [3]. The improved QuickSort algorithm can be applied on the sorting large-scale data, and exhibits slightly superior computational efficiency compared to classical sequential QuickSort. Jugé et al. proposed an adaptive ShiversSort algorithm for efficiently sorting partially sorted data, which is considered as a variant of the well-known algorithm TimSort [13]. Yang et al. proposed a high-performance parallel sorting algorithm on a CPU-DSP heterogeneous processor [27]. These methods typically take into account different sorting settings, such as parallel sorting environments, data criterions, or specific CPU architectures. These large-scale sorting algorithms provide an efficient post-processing step for collecting the statistics of large-scale IP address data.
@@ 2.3 Parallel hardware architectures
The hardware architecture of modern processors usually consists of more than two independent central processing units (CPUs) or graphics processing units (GPUs). Parallel software platforms can be implemented using high-level programming frameworks for specific hardware architectures [5]. The Compute Unified Device Architecture (CUDA) is a parallel computing platform for general computing on GPUs. Most parallel sorting algorithms are variants of standard, well-known sorting algorithms adapted to GPU hardware architecture. For example, Cederman designed a quick sort for the GPU platform [4], and Peters proposed an adaptive bitonic sorting algorithm with a bitonic tree for GPUs [20]. The parallel computation of sorting algorithms is considered to be the most efficient way of sorting elements on parallel hardware architectures [26].
@@ 3.1 Problem Formulation
IP flow data $FD$ is a sequence of IP records, that is, $FD=\left\{\left({x_{1},p_{1}}\right),...,\left({x_{n},p_{n}}\right)\right\}$ and ${n\geq 1{e^{6}}}$, where each pair of elements $\left({x_{i},p_{i}}\right)$ $\left({i\in\left[{1,n}\right]}\right)$ consists of an IP address $x_{i}$ and a set of corresponding user behavior attributes $p_{i}$. Given a finite set of IP addresses $X=\left\{{{x_{1}},{x_{2}},...,{x_{n}}}\right\}\in{\mathbb{R}^{m\times n}}$, the purpose of the IP address statistics task is to efficiently determine the first $k$ IP addresses of the most frequent occurrences in $X$, where $m$ the dimensionality of an individual IP address and $k\ll n$.
@@ 3.3.1 First method: TLMB
We formally present a storage strategy for IP addresses that consists of two layers that consist of a limited number of memory blocks. The first layer contains $256\times 256$ memory blocks. The first three parts of the IP address can be mapped into the corresponding position of the element in a particular memory block of the first layer according to the individual values of the three parts. We allocate a memory block in the other layer for the IP address when its first three parts are initially given. Each element of a memory block in this layer stores the number of occurrences of the corresponding IP address. Figure 2 shows an example of the relationship mapping between the memory blocks of two layers and an IP addresses.
Consider the individual IP address 1.2.3.4, we have $1\times 255\times 255+2\times 255=65535$, which represents the index of the memory block in the first layer. The positions of the first two dimensions of the sparse matrix can be mapped to the elements of the memory blocks included in the first layer. The third part of the IP address denotes the index of the memory block in the second layer. The positions of the final dimensionality of the sparse matrix can be indexed by combining the starting address of the memory block in the second layer with the integer value of the four parts of the IP address.
@@ 4.1 Experimental Settings
In this section, we evaluate the performance of the proposed methods  [^2] https://github.com/chenjie20/IPStatistics on two different types of datasets, i.e., three synthetic datasets and two real-world datasets. The three synthetic datasets contain 5 million, 10 million, and 50 million randomly generated IP records. Each individual IP address contains one or more of IP records. The average number of IP records is 100 for each individual IP address. In particular, the IPv4 addresses were specifically divided into four segments in the experiments. These experimental settings ensure a comprehensive evaluation of the capacity of TLMB and SSMB to efficiently collect statistics for large-scale IP address data. Parameter $k$ represents the number of frequently occurring IP addresses. Additionally, two real-world network traffic datasets, provided by the Center for Applied Internet Data Analysis (CAIDA)  [^3] https://www.caida.org/catalog/datasets/ipv4_prefix_probing_dataset, are collected from various parts of the internet. These two datasets are widely used in networking and traffic analysis research. The statistics of these datasets are summarized in Table 2.
Table 2: Statistics of the datasets.
Data sets IP Records Individual IP Addresses Size Type
1 5, 000, 000 50, 000 77.5 MB Synthetic
2 10, 000, 000 100, 000 155 MB Synthetic
3 50, 000, 000 500, 000 775 MB Synthetic
4 1, 114, 633 107, 988 14.4 MB Real
5 1, 430, 258 133, 116 18.5 MB Real
@@ 2. $\bullet$
Memory blocks were pre-allocated for TLMB and SSMB, with sizes ranging from small-scale (e.g., 5 million) to large-scale (e.g., 50 million) to test scalability. Two metrics were employed to evaluate the sorting performance, that is, computational cost and memory use. All experiments were implemented using the C language on a Windows platform with an Intel i7-9700k CPU and 32 GB RAM.
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
1 10 1,458.14 (2.32) 15.18 (0.07) 2.51 (0.01) 18.43 (0.05)
100 1,458.56 (2.76) 15.21 (0.04) 2.52 (0.01) 18.43 (0.06)
2 10 2,927.23 (11.64) 17.41 (0.11) 4.92 (0.01) 27.59 (0.05)
100 2,934.65 (28.93) 17.45 (0.05) 4.95 (0.01) 27.65 (0.06)
3 10 14,547.09 (13.95) 35.33 (0.09) 24.22 (0.07) 101.01 (0.17)
100 14,566.36 (26.78) 35.39 (0.05) 24.24 (0.04) 101.20 (0.2)
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
1 10 43.63 (0.12) 16,332.81 (0.07) 189.61 (0.16) 139.77 (0.13)
100 43.73 (0.06) 16,332.77 (0.05) 189.64 (0.07) 139.79 (0.07)
2 10 74.8 (0.12) 16,332.75 (0.12) 239.11 (0.1) 139.74 (0.13)
100 74.83 (0.05) 16,332.77 (0.07) 239.06 (0.15) 139.7 (0.12)
3 10 324.45 (0.99) 16,332.59 (0.06) 628.33 (0.07) 139.6 (0.15)
100 324.4 (1.02) 16,332.6 (0.16) 628.27 (0.01) 139.71 (0.16)
@@ 4.2.1 Experimental Evaluation on Synthesized Data
Table 5: Computational cost (s) of different methods on the two real-world datasets.
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
4 10 1675.3 6.73 0.61 7.00
100 1673.9 6.90 0.62 7.09
5 10 1996.8 7.33 0.78 6.88
100 1989.9 7.47 0.80 7.02
Table 6: Memory use (MB) of different methods on the two real-world datasets.
Data $k$ Hash Mapping IP Mapping Ours (TLMB) Ours (SSMB)
4 10 238.9 16,392.1 249.8 183.3
100 238.8 16,392.1 249.8 183.4
5 10 266.7 16,391.9 281.4 183.4
100 266.7 16,392.3 281.4 183.4
4.2.2 Experimental Evaluation on Real-World Data
We evaluate the proposed and competing methods on two real-world datasets. Tables 5 and 6 show the computational costs and memory usages levels of different methods, respectively. TLMB consistently incurs lower computational costs than do the other methods. SSMB and IP mapping demonstrate comparable computational costs. However, SSMB significantly reduces the memory requirements relative to IP mapping. Furthermore, Hash mapping incurs a higher computational cost than the competing methods do because of its use of hash computations, making it prohibitively time-consuming in practice. As the number of IP addresses increases across the two real-world datasets, SSMB maintains relatively stable memory usage. These findings highlight the effectiveness of both TLMB and SSMB.
4.3 Ablation study
To investigate the impact of the memory blocks in the proposed TLMB and SSMB methods, we performed ablation studies on the three synthetic datasets. Specially, we examined two particular cases in the experiments. The hash table was employed to replace the second part of TLMB and SSMB, respectively. The primary goal of the ablation study is to demonstrate the importance of the memory blocks on collecting the statistics of large-scale IP address data. The invariants of TLMB and SSMB corresponding to these two cases are referred to as TLMBhash and SSMBhash, respectively.
Tables 7 and 8 show the results of the ablation study regarding the computational cost and memory use. TLMBhash exhibits similar computational cost and memory use compared to SSMBhash on the first two synthetic datasets. Additionally, the computational cost of TLMBhash is slightly higher than that of SSMBhash, while its memory usage is marginally lower. TLMBhash and SSMBhash achieve an acceptable computational cost. However, TLMB and SSMB achieves superior performance in computational cost and memory use compared with those of TLMBhash and SSMBhash. These results further emphasize that integrating a hash table scheme into TLMB and SSMB is both time-consuming and memory-intensive. Therefore, the results of the ablation study demonstrate the effectiveness of the memory blocks in the proposed TLMB and SSMB methods.
Table 7: Ablation study concerning the computational costs (s) incurred on the three synthetic datasets.
Data $k$ TLMBhash SSMBhash Ours (TLMB) Ours (SSMB)
1 10 37.95 37.16 2.51 18.43
100 37.9 37.34 2.52 18.43
2 10 75.68 74.51 4.92 27.59
100 75.19 74.05 4.95 27.65
3 10 384.72 374.49 24.22 101.01
100 384.23 375.9 24.24 101.20
Table 8: Ablation study concerning the memory use (MB) required for the three synthetic datasets.
Data $k$ TLMBhash SSMBhash Ours (TLMB) Ours (SSMB)
1 10 3,397.3 3397.2 189.61 139.77
100 3,397.5 3,397.2 189.64 139.79
2 10 6,654.4 6,654.3 239.11 139.74
100 6,654.4 6,654.4 239.06 139.7
3 10 25,665.6 27,367.6 628.33 139.6
100 26,008.6 26,517.5 628.27 139.71
4.4 Empirical investigation
We empirically examined the effect induced by varying the $k$ considered in the proposed TLMB and SSMB methods. Here $k$ was selected from the set $\left\{{10,20,50,100,200,500}\right\}$. The computational cost and memory use were employed to evaluate TLMB and SSMB with different $k$ values.
Fig. 3 shows the computational costs of TLMB and SSMB with different $k$ values. As expected, the computational cost gradually increases as the number of IP records increases from 5 million to 50 million. Moreover, the computational costs of TLMB and SSMB remain relatively stable when $k$ varies from 10 to 500 on each synthetic dataset. This finding demonstrates the stability of TLMB and SSMB for computational efficiency when collecting the statistics of large-scale IP address data. Fig. 4 shows the memory uses of TLMB and SSMB with different $k$ values. We observe that TLMB requires more memory use as the number of IP records grows. In contrast, SSMB maintains relatively stable memory across varying numbers of IP records. This finding indicates that SSMB can satisfy certain memory requirements when handling varying numbers of IP records.
Figure 3: The computational costs of the proposed TLMB and SSMB methods with different $k$ values. (a) TLMB and (b) SSMB.
Figure 4: The memory uses of the proposed TLMB and SSMB methods with different $k$ values. (a) TLMB and (b) SSMB.
4.5 Discussion
The gap in computational cost between TLMB and Hash Mapping dramatically increases when the number of IP records increases from 5 million to 50 million. This is because TLMB avoids hash collisions when an IP address is mapped to the corresponding memory block. The computational cost of TLMB is linearly proportional to the number of IP addresses. Moreover, the memory use of SSMB remains almost unchanged regardless of the number of IP addresses. This is consistent with the theory underlying the second proposed mapping mechanism. There is a negligible effect on the computational costs of TLMB and SSMB when $k$ increases from 10 to 100. Moreover, the changes in the memory use of TLMB and computational cost of SSMB are tolerable in practical applications as the number of IP addresses increases. Consequently, TLMB and SSMB reach a reasonable balance between computational cost and memory use when compared with Hash Mapping and IP Mapping. This reveals that the two relationship mapping mechanisms for memory blocks and IP addresses are effective approaches for the design of the four-dimensional sparse matrix.
The memory blocks designed in TLMB and SSMB exhibit superior relationship mapping capabilities compared to those of hash mapping. The memory block takes fully advantages of the inherent property of the memory address, which is employed to corresponding to each part of an IP address. This indicates that integrating the memory block into TLMB and SSMB is both time-stable and memory-stable. In contrast, hash mapping uses a pair of key and value to store the statistics of IP address data. Unfortunately, IP addresses are often sparse in practical applications. Hash mapping requires additional memory to store the remaining two or three parts of the IP address as keys corresponding to TLMB and SSMB, respectively. This has a significant negative impact on the memory use of hash mapping. Therefore, the proposed memory block significantly enhances the capacity of TLMB and SSMB in collecting the statistics of large-scale IP address data.